Three identical charges each of 1 micro coulomb are placed at the corners of an equilateral triangle of side 10 cm. Calculate the electric field intensity at the geometric centre of the triangle.
 
[I know that the answer is zero, but as per our school exams we have to find the value of electric field intensity due to every charge. SO PLEASE SOLVE BY FINDING THE VALUE OF ELECTRIC FIELD INTENSITY DUE TO EACH CHARGE !!!!!!]

Grade:12

1 Answers

venkat
6 years ago
Distance of each charge from the geometric centre of the equilateral triangle d=10/\sqrt{3}
Electric field due to each charge is given by E=1/q/(4\pi\epsilon _{0})d^{2}
Using the above formula find the electric field intensity due to each charge and add them vectorially to get the answer.